Ex. 4.1. In R, define a ◦ b = a + b − ab. Show that this binary operation is associative, and that (R, ◦) is a monoid with zero as the identity element.
1
Expert's answer
2012-10-25T11:13:30-0400
a ◦ b belongs to R
associative because
a ◦ (b ◦ c)=a ◦
(b+c-bc)=a+b+c-bc-a*(b+c-bc)=a+b+c-ab-ac-bc+abc
(a ◦ b) ◦ c = (a+b-ab) ◦
c=a+b-ab+c-(a+b-ab)c=a+b+c-ab-ac-bc+abc

zero is 0
a ◦
0=a+0-a*0=a
0 ◦ a=0+a-0*a=a
so R with operation ◦ is a monoid